# Finding product sets in some classes of amenable groups

### v1

@@ 4.2. A continuous ergodic decomposition
In this subsection we will define measures $\lambda_{(x_{1},x_{2})}$, for $(x_{1},x_{2})\in X\times X$, in a way that $(x_{1},x_{2})\mapsto\lambda_{(x_{1},x_{2})}$ will be a continuous ergodic decomposition of $\mu\times\mu$. We follow the definition given in [KMRR22, Eq. (3.10)] and [KMRR23b, Eq. (3.1)].
@@
For the proof of (i) and (ii) we refer to [KMRR22, Proposition 3.11], as the proof there can be directly adapted to our case. We will now prove (iii). It is not too difficult to see that for all $(x_{1},x_{2})\in X\times X$ and for all $g\in G$, $(T_{g}\times T_{g})\lambda_{(x_{1},x_{2})}=\lambda_{(x_{1},x_{2})}$, i.e., $\lambda_{(x_{1},x_{2})}$ is $T\times T$-invariant. Therefore, to prove (iii), it suffices to prove that there is some Følner sequence $\Psi$ in $G$ such that for $(\mu\times\mu)$-almost every $(x_{1},x_{2})\in X\times X$ and all bounded and measurable functions $F$ on $X\times X$,
@@
Step 2. Consider the sequence of probability measures on $K$ defined by $\nu_{N}:=\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}\delta_{\alpha(g)}$. From Lemma 4.3 we know that $\nu_{N}\to m_{K}$ as $N\to\infty$ in the ${}^{\ast}$ topology.
@@
We will need the following two lemmas, the proofs of which are omitted, as they can be found in [KMRR23b]:
@@
[KMRR23b, Lemma 3.8] Let $W$ be a compact metric space, $M(W)$ the space of Borel probability measures on $W$ endowed with the ${}^{\ast}$ topology, and $\mathcal{F}(W)$ the space of closed, non-empty subsets of $W$ with the Hausdorff metric. Then
@@
[KMRR23b, Lemma 3.9] The disintegration $z\mapsto\eta_{z}$ satisfies that the subset
@@
Using those, we can now prove the following proposition, which is a variant of [KMRR23b, Proposition 3.10].
@@
Let $F:K\to\mathcal{F}(X)$ given by $F(k)=\texttt{{supp}}(\eta_{kH})$. The natural projection $p:K\to Z$ is continuous, hence Borel measurable, the map $z\mapsto\eta_{z}$ is Borel measurable, as $(\eta_{z})_{z\in Z}$ is a disintegration and by Lemma 4.11, the map $\nu\mapsto\texttt{{supp}}(\nu)$ is also Borel measurable, thus, we obtain that their composition $F$ is also Borel measurable. By Lusin’s theorem [AB06, Theorem 12.8], for any $j\in\mathbb{N}$ there exists a closed $K_{j}\subset K$ with
@@
Let $S=\{(x_{1},x_{2})\in X\times X\colon(x_{1},x_{2})\in\texttt{{supp}}(\lambda_{(x_{1},x_{2})})\}$. By Lemma 4.11, $S$ is a Borel subset of $X\times X$. Consider a sequence $\delta_{j}\to 0$ such that Proposition 4.13 is satisfied, and let $L\subset X$ be the set of $x\in X$ that satisfy (4.11). By Proposition 4.13, $\mu(L)=1$. Following the argument in [KMRR23b, Proposition 3.11] we will show that $\sigma(L\times L)=1$ and $L\times L\subset S$. Consequently, we will have that $\sigma(S)=1$, concluding the proof.
@@
$=\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cdot k_{2}^{-1}k_{1}\cap\texttt{{B}}_{K}(k_{2},\delta)\cdot k_{2}^{-1}k_{1})}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta)\cdot k_{2}^{-1}k_{1})}$
$=\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cap\texttt{{B}}_{K}(k_{2},\delta))}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta))}$
(4.19) $\geq\frac{3}{4}.$
@@
It remains to show Theorem 4.10. To this end, we need the following lemma, which is the analog of [KMRR23b, Lemma 3.7] in our setting.
@@
In this proof we follow the argument used in the proof of [KMRR22, Theorem 7.6]. Apply Lemma 2.8 for the ergodic decomposition $(x_{0},x_{1})\mapsto\lambda_{(x_{0},x_{1})}$, to obtain a Følner sequence $\Phi$ such that
@@
where $\Psi$ is the Følner sequence defined by $\Psi_{k}=h_{k}\Phi_{N(k)}$, for $k\in\mathbb{N}$. Using (4.6), the above equation becomes
@@
We start by showing Theorem 1.7, and then we will prove Corollaries 1.8 and 1.9. We split the proof of Theorem 1.7 into the following three lemmas:
@@
Let $G,\phi,\Phi,\Psi$ and $\eta$ be as in the assumptions of Lemma 5.1. We will prove something stronger than we require, namely that for any $u:G\to[0,1]$ we have
@@
and then for any $\varepsilon>0$, taking $\delta=\eta\varepsilon/M>0$ yields that $G$ is $\phi$-absolutely continuous. Let $u:G\to[0,1]$, and let $(\Psi_{N_{k}})_{k\in\mathbb{N}}$ such that
@@
Let $G$ be a torsion-free finitely generated nilpotent group and fix a Mal’cev coordinate system $(t_{1},\dots,t_{s})$. By [KM79, Theorem 17.2.5], we have that for any $1\leq i\leq s$, there are polynomials $p_{i}$ in $i-1$ variables, with rational coefficients, satisfying $p_{i}(\mathbb{Z}^{i-1})\subset\mathbb{Z}$ such that for any $x\in G$, we have
(5.2) $t_{i}(x^{2})=2t_{i}(x)+p_{i}(\{t_{j}(x)\colon j<i\}).$
@@
The proof becomes obvious by using (5.2). ∎
@@
It is not hard to check that $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is Følner sequence in $G$. We show that $s(\Psi_{N})\subset\Psi_{N+1}$ for every $N\in\mathbb{N}$. Let $N\in\mathbb{N}$ and then
@@
Let $x^{2}=(t_{1}(x^{2}),\dots,t_{s}(x^{2}))\in s(\Psi_{N})$. We claim that for each $1\leq i\leq s$, we have that
@@
where the last one follows from that $|p_{i}((t_{j}(x))_{1\leq j<i})|\leq\gamma_{i}b_{i}c_{i}^{(N-1)d_{i}}$, by (5.3), and from the definition of $c_{i}$ along with that $d_{i}\geq 1$. This shows (5.4) and hence that $s(\Psi_{N})\subset\Psi_{N+1}$.
It remains to prove (5.1) and we show it for the Følner $\Psi$ and with $\eta:=\prod_{1\leq i\leq s}c_{i}^{-d_{i}}>0$. In other words, we prove that for any $N\in\mathbb{N}$ we have
(5.5) $\frac{|s(\Psi_{N})|}{|\Psi_{N+1}|}=\prod_{1\leq i\leq s}c_{i}^{-d_{i}}.$
@@
$|s(\Psi_{N})|=|\Psi_{N}|=\prod_{1\leq i\leq s}2^{s}\bigg{(}\prod_{1\leq i\leq s}c_{i}^{d_{i}}\bigg{)}^{N-1}.$
@@
$|\Psi_{N+1}|=\prod_{1\leq i\leq s}2^{s}\bigg{(}\prod_{1\leq i\leq s}c_{i}^{d_{i}}\bigg{)}^{N},$
@@
Let $s\in\mathbb{N}$ and $1\leq i\leq s$. We say that a non-empty set $L\subset\mathbb{Z}^{s}$ is a line in the $i$-th coordinate if there is a non-empty interval $I$ in $\mathbb{Z}$, i.e. a set of the form $\{m,\dots,m+r\}$ for some $m,r\in\mathbb{Z}$ with $r\geq 0$, and some integers $x_{j}$, $j\in\{1,\dots,s\}\setminus\{i\}$ such that
@@
We refer to the length of the interval $I$ as the length of the line $L$.
@@
The first part of Corollary 1.8 follows immediately by combining Theorems 1.5 and 1.7. It remains to prove that if $G$ is a torsion-free finitely generated nilpotent group, then given a Mal’cev coordinate system $(t_{1},\dots,t_{s})$ on $G$, we can choose $B$ so that for any finite set $C\subset\mathbb{Z}$ and any $1\leq i\leq s$, the set $\{b\in B:t_{i}(b)\in C\}$ is finite.
@@
Let $G$ be an amenable group, $H$ be a subgroup of $G$ with $|G:H|=r<\infty$ and $\Phi$ be a Følner sequence in $G$. Then the following hold:
@@
Let $G$ be a finitely generated virtually nilpotent group, $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ be a Følner sequence in $G$ and $A$ be a subset of $G$ with $\overline{d}_{\Phi}(A)>0$. By passing to a subsequence for which the limit exists, we may assume that $d_{\Phi}(A)>0$. Let $G^{\prime}$ be a nilpotent finite-index subgroup of $G$. Since $G$ is finitely generated, by writing it as a disjoint union of finitely many left cosets of $G^{\prime}$, it is easy to see that $G^{\prime}$ is also finitely generated, and it is also nilpotent. By [KM79, Theorem 17.2.2], there exists a normal subgroup $H$ of $G^{\prime}$ with finite index, which is torsion-free. Hence, $H$ is a torsion-free finitely generated nilpotent group, which has finite index in $G$. By writing $G$ as finite disjoint union of left cosets of $H$, we can see that there exists some $g\in G$ such that $\overline{d}_{\Phi}(g^{-1}A\cap H)=\overline{d}_{\Phi}(A\cap gH)>0$. Again, by passing to a subsequence, we may assume that the limit exists, so $d_{\Phi}(g^{-1}A\cap H)=d_{\Phi}(A\cap gH)>0$. We let $\Psi_{N}=\Phi_{N}\cap H$ for each $N\in\mathbb{N}$ and by Lemma 5.4 (ii), $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is a Følner sequence on $H$. Hence for every $N\in\mathbb{N}$ we have that
@@
Once we have established the corollaries concerning finitely generated nilpotent groups, we move on to showing the results corresponding to abelian groups, namely Theorem 1.11 and Corollary 1.12. Obviously, the latter is an immediate consequence of the former and of Theorem 1.5, so it suffices to show Theorem 1.11.
@@
Let $(G,+)$ be an abelian group such that $2G$ has finite index in $G$, let $\Psi$ be any Følner sequence in $G$ and consider the doubling map $s_{G}:G\to 2G,g\mapsto 2g$. We will show that the assumptions of Lemma 5.1 are satisfied, which will yield that $G$ is square absolutely continuous. Then $|\ker(s_{G})|=|G:2G|=r$, hence
(5.6) $|s_{G}^{-1}(\{g\})|\leq r,\quad\forall g\in G.$
By assumption, there exist $r\in\mathbb{N}$ and $g_{1},\dots,g_{r}\in G$, where $g_{1}=e_{G}$, such that
$G=\bigsqcup_{i=1}^{r}(g_{i}+2G).$
We define $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ by $\Phi_{N}=\bigsqcup_{i=1}^{r}(g_{i}+2\Psi_{N})$. Assuming that this is a Følner sequence in $G$, by taking $\eta:=\frac{1}{r}$, we have that
(5.7) $s_{G}(\Psi_{N})=2\Psi_{N}\subset\Phi_{N}\qquad\text{and}\qquad\frac{|s_{G}(\Psi_{N})|}{|\Phi_{N}|}=\frac{|2\Psi_{N}|}{r|2\Psi_{N}|}=\eta.$
Combining (5.6) and (5.7), we have that the assumptions of Lemma 5.1 are indeed satisfied. Thus, it suffices to show that $\Phi$ is a Følner sequence in $G$.
Let $g\in G$ and $\varepsilon>0$. There exist $1\leq i_{0}\leq r$ and $h\in G$ such that $g=g_{i_{0}}+2h$. Then
(5.8) $(g+\Phi_{N})\cap\Phi_{N}=\bigg{(}\bigsqcup_{i=1}^{r}(g_{i_{0}}+g_{i}+2(h+\Psi_{N}))\bigg{)}\cap\bigg{(}\bigsqcup_{j=1}^{r}(g_{j}+2\Psi_{N})\bigg{)}.$
Since the cosets $g_{j}+2G$ are disjoint, it follows that for each $1\leq i\leq r$, there exists a unique $1\leq j(i)\leq r$ such that $g_{i_{0}}+g_{i}\in g_{j(i)}+2G$, so there is $y_{i}\in G$ such that $g_{i_{0}}+g_{i}=g_{j(i)}+2y_{i}$. In addition, if we assume that for $i_{1}\neq i_{2}$ we have $j(i_{1})=j(i_{2})$, then we have that $g_{i_{1}}-g_{i_{2}}=2y_{i_{1}}-2y_{i_{2}}\in 2G$, so $g_{i_{1}}+2G=g_{i_{2}}+2G$, which is a contradiction. Therefore, the map $j:\{1,\dots,r\}\to\{1,\dots,r\},i\mapsto j(i)$ is a bijection, and then (5.8) becomes
$(g+\Phi_{N})\cap\Phi_{N}=\bigsqcup_{i=1}^{r}\big{(}(g_{j(i)}+2(y_{i}+h+\Psi_{N}))\cap(g_{j(i)}+2\Psi_{N})\big{)}=\bigsqcup_{i=1}^{r}\big{(}g_{j(i)}+(2(y_{i}+h+\Psi_{N})\cap 2\Psi_{N})),$
@@
Now since $\Psi$ is a Følner sequence in $G$, for $N$ sufficiently large, we have that for every $1\leq i\leq r$, $|(y_{i}+h+\Psi_{N})\triangle\Psi_{N}|\leq\frac{\varepsilon}{r}|\Psi_{N}|$, and then we have that
(5.10) $\frac{|s_{G}((y_{i}+h+\Psi_{N})\cap\Psi_{N})|}{|s_{G}(\Psi_{N})|}\geq 1-\frac{|s_{G}(\Psi_{N}\triangle(y_{i}+h+\Psi_{N}))|}{|s_{G}(\Psi_{N})|}\geq 1-\frac{r|\Psi_{N}\triangle(y_{i}+h+\Psi_{N})|}{|\Psi_{N}|}\geq 1-\varepsilon,$
where for the second inequality above we use that $|s_{G}(\Psi_{N}\triangle(y_{j}+h+\Psi_{N}))|\leq|(\Psi_{N}\triangle(y_{j}+h+\Psi_{N}))|$ and that $|s_{G}(\Psi_{N})|\geq\frac{1}{r}|\Psi_{N}|$, as $|s_{G}^{-1}(\{z\})|\leq r$ for all $z\in G$.
Then, combining (5) and (5.10) we get that for $N$ sufficiently large
$\frac{|(g+\Phi_{N})\cap\Phi_{N}|}{|\Phi_{N}|}\geq\frac{\sum_{i=1}^{r}(1-\varepsilon)|s_{G}(\Psi_{N})|}{r|s_{G}(\Psi_{N})|}=1-\varepsilon.$

### v3

@@ 4.2. A continuous ergodic decomposition
In this subsection we will define measures $\lambda_{(x_{1},x_{2})}$, for $(x_{1},x_{2})\in X\times X$, in a way that $(x_{1},x_{2})\mapsto\lambda_{(x_{1},x_{2})}$ will be a continuous ergodic decomposition of $\mu\times\mu$, i.e., $(x_{1},x_{2})\mapsto\lambda_{(x_{1},x_{2})}$ will be both a continuous map and an ergodic decomposition of $\mu\times\mu$. We follow the definition given in [14, Eq. (3.10)] and [15, Eq. (3.1)].
@@
For the proof of (i) and (ii) we refer to [14, Proposition 3.11], as the proof there can be directly adapted to our case. We will now prove (iii). It is not too difficult to see that for all $(x_{1},x_{2})\in X\times X$ and for all $g\in G$, $(T_{g}\times T_{g})\lambda_{(x_{1},x_{2})}=\lambda_{(x_{1},x_{2})}$, i.e., $\lambda_{(x_{1},x_{2})}$ is $T\times T$-invariant. Therefore, to prove (iii), it suffices to prove that there is some Følner sequence $\Psi$ in $G$ such that for $(\mu\times\mu)$-almost every $(x_{1},x_{2})\in X\times X$ and all bounded and measurable functions $F$ on $X\times X$,
@@
Step 2. Consider the sequence of probability measures on $K$ defined by $\nu_{N}:=\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}\delta_{\alpha(g)}$. From Lemma 4.3 we know that $\nu_{N}\to m_{K}$ as $N\to\infty$ in the weak topology.
@@
We will need the following two lemmas, the proofs of which are omitted, as they can be found in [15]:
@@
[15, Lemma 3.8] Let $W$ be a compact metric space, $M(W)$ the space of Borel probability measures on $W$ endowed with the weak topology, and $\mathcal{F}(W)$ the space of closed, non-empty subsets of $W$ with the Hausdorff metric. Then
@@
[15, Lemma 3.9] The disintegration $z\mapsto\eta_{z}$ satisfies that the subset
@@
Using those, we can now prove the following proposition, which is a variant of [15, Proposition 3.10].
@@
Let $F:K\to\mathcal{F}(X)$ given by $F(k)=\texttt{{supp}}(\eta_{kH})$. The natural projection $p:K\to Z$ is continuous, hence Borel measurable, the map $z\mapsto\eta_{z}$ is Borel measurable, as $(\eta_{z})_{z\in Z}$ is a disintegration and by Lemma 4.11, the map $\nu\mapsto\texttt{{supp}}(\nu)$ is also Borel measurable, thus, we obtain that their composition $F$ is also Borel measurable. By Lusin’s theorem [2, Theorem 12.8], for any $j\in\mathbb{N}$ there exists a closed $K_{j}\subset K$ with
@@
Let $S=\{(x_{1},x_{2})\in X\times X\colon(x_{1},x_{2})\in\texttt{{supp}}(\lambda_{(x_{1},x_{2})})\}$. By Lemma 4.11, $S$ is a Borel subset of $X\times X$. Consider a sequence $\delta_{j}\to 0$ such that Proposition 4.13 is satisfied, and let $L\subset X$ be the set of $x\in X$ that satisfy (4.11). By Proposition 4.13, $\mu(L)=1$. Following the argument in [15, Proposition 3.11] we will show that $\sigma(L\times L)=1$ and $L\times L\subset S$. Consequently, we will have that $\sigma(S)=1$, concluding the proof.
@@
$\hskip 128.0374pt=\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cdot k_{2}^{-1}k_{1}\cap\texttt{{B}}_{K}(k_{2},\delta)\cdot k_{2}^{-1}k_{1})}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta)\cdot k_{2}^{-1}k_{1})}$
$\hskip 128.0374pt=\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cap\texttt{{B}}_{K}(k_{2},\delta))}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta))}$
(4.19) $\hskip 128.0374pt\geq\frac{3}{4}.$
@@
It remains to show Theorem 4.10. To this end, we need the following lemma, which is the analog of [15, Lemma 3.7] in our setting.
@@
In this proof we follow the argument used in the proof of [14, Theorem 7.6]. Apply Lemma 2.8 for the ergodic decomposition $(x_{0},x_{1})\mapsto\lambda_{(x_{0},x_{1})}$, to obtain a Følner sequence $\Phi$ such that
@@
where $\Psi$ is the Følner sequence defined by $\Psi_{k}=\Phi_{N(k)}h_{k}$, for $k\in\mathbb{N}$. Using (4.6), the above equation becomes
@@
We start by showing Theorem 1.8, and then we will prove Corollaries 1.9 and 1.10. We split the proof of Theorem 1.8 into the following three lemmas:
@@
Let $G,M,\phi,\Phi,\Psi$ and $\eta$ be as in the assumptions of Lemma 5.1. We will prove something stronger than we require, namely that for any $u:G\to[0,1]$ we have
@@
and then for any $\varepsilon>0$, taking $\delta=\eta\varepsilon/M>0$ yields that $G$ is $\phi$-absolutely continuous. Let $u:G\to[0,1]$, and let $(\Psi_{N_{k}})_{k\in\mathbb{N}}$ satisfy
@@
Let $G$ be a torsion-free finitely generated nilpotent group and fix a Mal’cev coordinate system $(t_{1},\dots,t_{s})$. By [11, Theorem 17.2.5], we have that for any $1\leq i\leq s$, there are polynomials $p_{i}$ in $i-1$ variables, with rational coefficients, satisfying $p_{i}(\mathbb{Z}^{i-1})\subset\mathbb{Z}$ such that for any $x\in G$, we have
(5.2) $t_{i}(x^{2})=2t_{i}(x)+p_{i}(t_{1}(x),\dots,t_{i-1}(x)).$
In the previous, $p_{1}$ is a polynomial in 00 variables, which means that it is a constant polynomial.
@@
The proof becomes obvious by using (5.2) and the injectivity of the coordinate map $G\to\mathbb{Z}^{s}$, $x\mapsto(t_{1}(x),\dots,t_{s}(x))$, which follows from the fact that $G$ is torsion-free. ∎
@@
It is not hard to check that $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is Følner sequence in $G$. We show that $s_{G}(\Psi_{N})\subset\Psi_{N+1}$ for every $N\in\mathbb{N}$. Let $N\in\mathbb{N}$ and then
@@
Let $x^{2}=(t_{1}(x^{2}),\dots,t_{s}(x^{2}))\in s_{G}(\Psi_{N})$. We claim that for each $1\leq i\leq s$, we have that
@@
where the last one follows from $|p_{i}((t_{j}(x))_{1\leq j<i})|\leq\gamma_{i}b_{i}c_{i}^{(N-1)d_{i}}$, by (5.3), and from the definition of $c_{i}$ along with that $d_{i}\geq 1$. This shows (5.4) and hence that ${s_{G}}(\Psi_{N})\subset\Psi_{N+1}$.
It remains to prove (5.1) and we show it for the Følner sequence $\Psi$ and with $\eta:=\prod_{1\leq i\leq s}c_{i}^{-d_{i}}>0$. In other words, we prove that for any $N\in\mathbb{N}$ we have
(5.5) $\frac{|{s_{G}}(\Psi_{N})|}{|\Psi_{N+1}|}=\prod_{1\leq i\leq s}c_{i}^{-d_{i}}.$
@@
$|s_{G}(\Psi_{N})|=|\Psi_{N}|=2^{s}\bigg{(}\prod_{1\leq i\leq s}c_{i}^{d_{i}}\bigg{)}^{N-1}.$
@@
$|\Psi_{N+1}|=2^{s}\bigg{(}\prod_{1\leq i\leq s}c_{i}^{d_{i}}\bigg{)}^{N},$
@@
Let $s\in\mathbb{N}$ and $1\leq i\leq s$. We say that a non-empty set $L\subset\mathbb{Z}^{s}$ is a line in the $i$-th coordinate if there is a non-empty interval $I$ in $\mathbb{Z}$, i.e., a set of the form $\{m,\dots,m+r\}$ for some $m,r\in\mathbb{Z}$ with $r\geq 0$, and some integers $x_{j}$, $j\in\{1,\dots,s\}\setminus\{i\}$ such that
@@
We refer to the cardinality of the interval $I$ as the length of the line $L$.
@@
The first part of Corollary 1.9 follows immediately by combining Theorems 1.6 and 1.8. It remains to prove that if $G$ is a torsion-free finitely generated nilpotent group, then given a Mal’cev coordinate system $(t_{1},\dots,t_{s})$ on $G$, we can choose $B$ so that for any finite set $C\subset\mathbb{Z}$ and any $1\leq i\leq s$, the set $\{b\in B:t_{i}(b)\in C\}$ is finite.
@@
Let $G$ be an amenable group, $H$ be a subgroup of $G$ with $[G:H]=r<\infty$ and $\Phi$ be a Følner sequence in $G$. Then the following hold:
@@
Let $G$ be a finitely generated virtually nilpotent group, $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ be a Følner sequence in $G$ and $A$ be a subset of $G$ with $\overline{d}_{\Phi}(A)>0$. By passing to a subsequence for which the limit exists, we may assume that $d_{\Phi}(A)>0$. Let $G^{\prime}$ be a nilpotent finite-index subgroup of $G$. Since $G$ is finitely generated, by writing it as a disjoint union of finitely many left cosets of $G^{\prime}$, it is easy to see that $G^{\prime}$ is also finitely generated, and it is also nilpotent. By [11, Theorem 17.2.2], there exists a normal subgroup $H$ of $G^{\prime}$ with finite index, which is torsion-free. Hence, $H$ is a torsion-free finitely generated nilpotent group, which has finite index in $G$. By writing $G$ as finite disjoint union of left cosets of $H$, we can see that there exists some $g\in G$ such that $\overline{d}_{\Phi}(g^{-1}A\cap H)=\overline{d}_{\Phi}(A\cap gH)>0$. Again, by passing to a subsequence, we may assume that the limit exists, so $d_{\Phi}(g^{-1}A\cap H)=d_{\Phi}(A\cap gH)>0$. We let $\Psi_{N}=\Phi_{N}\cap H$ for each $N\in\mathbb{N}$ and by Lemma 5.4 (ii), $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is a Følner sequence on $H$. Hence for every $N\in\mathbb{N}$ we have that
@@
Having established the corollaries concerning finitely generated nilpotent groups, we move on to showing the results corresponding to abelian groups, namely Theorem 1.12 and Corollary 1.13. Obviously, the latter is an immediate consequence of the former and of Theorem 1.6, so it suffices to show Theorem 1.12.
@@
[…]
For each $N$ we define $\Phi_{N}:=\bigsqcup_{i=1}^{r}\beta_{i}+\widetilde{\Phi}_{N}$ and we observe that $|\Phi_{N}|=r|\widetilde{\Phi}_{N}|$.
Claim $1$: $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ is a Følner in $G$.
Proof of Claim $1$.
Let $y\in G$ and $\varepsilon>0$. Then there exist $1\leq i_{0}\leq r$ and $h\in G$ such that $y=\beta_{i_{0}}+2h$. Then
(5.6) $(y+\Phi_{N})\cap\Phi_{N}=\bigg{(}\bigsqcup_{i=1}^{r}(\beta_{i_{0}}+\beta_{i}+2h+\widetilde{\Phi}_{N}))\bigg{)}\cap\bigg{(}\bigsqcup_{j=1}^{r}(\beta_{j}+\widetilde{\Phi}_{N})\bigg{)}.$
Since the cosets $\beta_{j}+2G$ are disjoint, it follows that for each $1\leq i\leq r$, there exists a unique $1\leq j(i)\leq r$ such that $\beta_{i_{0}}+\beta_{i}\in\beta_{j(i)}+2G$, so there is $h_{i}\in G$ such that $\beta_{i_{0}}+\beta_{i}=\beta_{j(i)}+2h_{i}$. In addition, if we assume that for $i_{1}\neq i_{2}$ we have $j(i_{1})=j(i_{2})$, then we have that $\beta_{i_{1}}-\beta_{i_{2}}=2h_{i_{1}}-2h_{i_{2}}\in 2G$, so $\beta_{i_{1}}+2G=\beta_{i_{2}}+2G$, which is a contradiction. Therefore, the map $j:\{1,\dots,r\}\to\{1,\dots,r\},i\mapsto j(i)$ is a bijection, and then (5.6) becomes
$(y+\Phi_{N})\cap\Phi_{N}=\bigsqcup_{i=1}^{r}\big{(}(\beta_{j(i)}+2h_{i}+2h+\widetilde{\Phi}_{N}))\cap(\beta_{j(i)}+\widetilde{\Phi}_{N})\big{)}=\bigsqcup_{i=1}^{r}\big{(}\beta_{j(i)}+((2h_{i}+2h+\widetilde{\Phi}_{N})\cap\widetilde{\Phi}_{N})),$
@@
Now since $\widetilde{\Phi}$ is a Følner sequence in $2G$, for $N$ sufficiently large we have that for every $1\leq i\leq r$, $|(2h_{i}+2h+\widetilde{\Phi}_{N})\triangle\widetilde{\Phi}_{N}|\leq\varepsilon|\widetilde{\Phi}_{N}|$, and then we have that
(5.8) $\frac{|(2h_{i}+2h+\widetilde{\Phi}_{N})\cap\widetilde{\Phi}_{N})|}{|\widetilde{\Phi}_{N}|}\geq 1-\frac{|(2h_{i}+2h+\widetilde{\Phi}_{N})\triangle\widetilde{\Phi}_{N}|}{|\widetilde{\Phi}_{N}|}\geq 1-\varepsilon.$
Then, combining (5.7) and (5.8) we get that for $N$ sufficiently large
$\frac{|(y+\Phi_{N})\cap\Phi_{N}|}{|\Phi_{N}|}=\frac{\sum_{i=1}^{r}|(2h_{i}+2h+\widetilde{\Phi}_{N})\cap\widetilde{\Phi}_{N})|}{r|\widetilde{\Phi}_{N}|}\geq\frac{\sum_{i=1}^{r}(1-\varepsilon)|\widetilde{\Phi}_{N}|}{r|\widetilde{\Phi}_{N}|}=1-\varepsilon.$